# Singular flat bands in three dimensions: Landau level spreading, quantum geometry, and Weyl reconstruction

### v1

@@ I Introduction
Electronic systems featuring flat bands, where the bandwidth is vanishingly small compared to other relevant energy scales, have long been of interest in condensed matter physics, driven by their unusual single-particle properties [1; 2; 3; 4; 5], strongly correlated phases [6; 7; 8; 9; 10; 11; 12; 13; 14; 15; 16; 17; 18; 19; 20; 21; 22], and experimental realizations [23; 24; 25; 26; 27; 28; 29; 30; 31; 32; 33; 34; 35; 36; 37; 38; 39; 40; 41; 42; 43; 44; 45]. A flat band can arise trivially when interatomic hopping is suppressed, leading to a so-called trivial flat band. In contrast, nontrivial flat bands emerge despite finite hopping amplitudes, as a result of destructive interference encoded in the lattice geometry. Prototypical examples include the dice [1], Lieb [7], and kagome [9] lattices. In such systems, the flat band is generally not isolated but instead touches a dispersive band at discrete points in momentum space. Considerable theoretical effort has been devoted to elucidating the mathematical origins and topological characteristics of these nontrivial flat bands [1; 8; 9; 10; 6; 46; 47; 2; 48; 49; 50; 51; 52; 53].
Recent studies have highlighted the role of nontrivial topology and quantum geometry in flat-band systems [54; 2; 55; 4; 56; 57; 58; 59; 60; 61; 60; 62], particularly in two-dimensional (2D) settings. In this context, nontrivial flat bands are often referred to as singular flat bands, as the eigenvector as a function of the Bloch momentum exhibits a singularity at a band touching point [4; 54; 5], where the quantum metric diverges. This singularity is not merely a mathematical feature; it also has direct physical consequences. A striking effect appears in the Landau level structure under an external magnetic field. In a trivial flat band, the electronic states remain degenerate in the presence of a magnetic field, as the localized electrons are insensitive to the Aharonov–Bohm phase. By contrast, in a singular flat band, the magnetic field induces Landau levels spreading over a finite energy range reflecting the underlying geometric singularity [4; 54].
@@
A key open question is whether singular flat bands can exist in three dimensions (3D) and, if so, how they respond to an external magnetic field. A promising candidate is the pyrochlore lattice, which hosts flat bands extending along all three momentum directions $(k_{x},k_{y},k_{z})$, with a band touching point at $\bm{k}=0$ [9; 2; 3]. This structure can be regarded as a 3D analog of the kagome lattice. Desipite its potential significance, the quantum geometric properties of these 3D flat bands, and their Landau level structure under magnetic fields, remain largely unexplored.
In this work, we investigate the geometric singularity of the 3D flat band in the pyrochlore lattice and examine its response to an external magnetic field. Our analysis is based on a three-orbital effective continuum model, which captures the electronic structure near the threefold-degenerate band-touching point. We show that the touching point is a point-like topological singularity in three-dimensional momentum space, as illustrated in Fig. 1(a). At any $k$-point, the eigenvectors can be represented by three-component real vectors. Notably, the flat band in the 3D system is twofold degenerate unlike in 2D cases, and hence its eigenvectors span a two-dimensional subspace within the three-component vector space, illustrated as green flat planes in Fig. 1(a). The presence of the point-like singularity is manifested in a hedgehog structure of a normal director perpendicular to the plane of the flat-band manifold. Additionally, the quantum geometric tensor for the flat bands acquires a non-Abelian structure  [63; 64; 65]: each of its component is a $2\times 2$ matrix, reflecting the intrinsic degeneracy of the flat bands.
@@
The three-dimensionality and intrinsic degeneracy introduce new features in the Landau level structure of the flat bands that are absent in 2D systems. For sufficiently large $|k_{z}|$, the degenerate flat bands are split into two branches by the orbital Zeeman effect, and the Landau levels within each branch remain nearly degenerate without energy spreading [Fig. 1(c)]. As $k_{z}$ decreases, the Landau levels in each branch acquire a finite energy spread, consistent with the increasing quantum metric. Near $k_{z}\sim 0$, hybridization between the two flat bands and the adjacent dispersive band gives rise to an intricate Landau level structure, potentially reflecting non-Abelian quantum geometry. We show that the full $k_{z}$-dependent behavior is well captured by an effective band model that incorporates the orbital Zeeman effect while neglecting full Landau quantization. Within this approximation, the system effectively behaves as a Weyl semimetal, and the Landau level spreading can be interpreted as a dispersion of the flat bands induced by orbital magnetism. The model also explains the presence of a chiral Landau level connecting the flat bands to the dispersive band — a characteristic feature of Weyl semimetals [66; 67; 68; 69; 70; 71].
@@
Lastly, while our effective continuum model for the pyrochlore lattice is constructed from three orbitals correponding to orbital angular momentum $\ell=1$, the framework can be naturally extended to arbtirary angular momentum quantum numbers $\ell$, described by $12\ell+$ 1 dimensional matrix. In particular, $\ell=3/2$ corresponds to Luttinger-Kohn-type model with parameters tuned to yield a flat band. While the original Luttinger-Kohn model with $\ell=3/2$ describes dispersive bands and has been studied in the context of various topological materials [72; 73; 74; 75; 76; 77; 78; 79; 80; 81; 82; 83; 84; 85; 86; 87; 88; 89; 90; 91; 92; 93; 94; 95; 96], our work focuses on a flat band variant that can be generalized to arbitrary $\ell$. The extension demonstrates that the essential features of the 3D singular flat band identified in pyrochlore lattice, such as Landau level spreading and emergent Weyl semimetal like behevior, are robust across a wide class of band models.
@@ II.1 Effective continuum model
We consider a pyrochlore lattice, as illustrated in Fig. 2(a). The primitive lattice vectors are defined as $\bm{a}_{1}=d(\hat{\bm{x}}+\hat{\bm{y}})$, $\bm{a}_{2}=d(\hat{\bm{y}}+\hat{\bm{z}})$, and $\bm{a}_{3}=d(\hat{\bm{z}}+\hat{\bm{x}})$, where $\hat{\bm{x}}$, $\hat{\bm{y}}$, and $\hat{\bm{z}}$ are unit vectors along $x$-, $y$-, and $z$-directions, respectively. The nearest neighbor intersite distance is $d/\sqrt{2}$. The unit cell consists of four sublattices with positions measured from the unit cell center [Fig. 2(a)],
@@
$(H_{\bm{k}})_{X,X^{\prime}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=}\langle\bm{k},\bm{\tau}_{X}|H|\bm{k},\bm{\tau}_{X^{\prime}}\rangle$
@@
$L_{x}=\left(\begin{array}[]{ccc}1&0&0\\ 0&0&-i\\ 0&i&0\\ \end{array}\right),\ L_{y}=\left(\begin{array}[]{ccc}0&0&i\\ 0&1&0\\ -i&0&0\\ \end{array}\right),$ (14)
$L_{z}=\left(\begin{array}[]{ccc}0&-i&0\\ i&0&0\\ 0&0&1\\ \end{array}\right),$ (18)
which act as generators of the rotation operators in Eq. (7). These $L_{i}$ satisfy the standard algebra $[L_{\alpha},L_{\beta}]=i\epsilon_{\alpha\beta\gamma}L_{\gamma}$, and represent effective orbital angular momentum associated with the sublattice degrees of freedom. Using them, the Hamiltonian in Eq. (6) can be rewritten as
@@ III.1 Landau level structure
We investigate the electronic structure of the 3D flat band system described by Eq. (19) under a uniform magnetic field $\bm{B}$ applied along the $z$-direction. We define the vector potential $\bm{A}$ satisfying $\bm{B}=\bm{\nabla}\times\bm{A}$. The Hamiltonian is then obtained by a standard process replacing the momentum components $k_{\alpha}$ in Eq. (19) with $\pi_{\alpha}/\hbar$ ($\alpha=x,y$), where $\bm{\pi}=-i\hbar\bm{\nabla}+e\bm{A}$ and the quadratic term is replaced with the symmetrized product $k_{\alpha}k_{\beta}\rightarrow(\pi_{\alpha}\pi_{\beta}+\pi_{\beta}\pi_{\alpha})/(2\hbar^{2})$. The momentum component along the magnetic field, $k_{z}$, remains a quantum number. We define the ladder operators $a=(\pi_{x}-i\pi_{y})/\sqrt{2\hbar eB}$ and $a^{\dagger}=(\pi_{x}+i\pi_{y})/\sqrt{2\hbar eB}$, which satisfy the commutation relation $[a,a^{\dagger}]=1$, noting that $[\pi_{x},\pi_{y}]=i\hbar eB$. As a result, the Hamiltonian in the magnetic field is written as
@@
Here, $\phi_{n}(x,y)$ denotes the wave function of $n$-th Landau levels in two-dimensional system with $n=0,1,2,\cdots$, where the index for the Landau level degeneracy in each $n_{\rm LL}$ is omitted. The $\vec{e}_{m}$ is a three-dimensional eigenvector of the orbital angular momentum $L_{z}$ with the eigenvalue $m=-1,0,1$. The basis of Eq. (26) satisfies the following equations,
@@
For $j=-1$ and 00, the eigenvectors reduce to one- and two-component forms, respectively, since the basis states $|k_{z},n,m\rangle$ with $n<0$ vanish. Correspondingly, the Hamiltonian matrices for $j=-1$ and $j=0$ are given by the $1\times 1$ and $2\times 2$ lower-right submatrices of Eq. (30), respectively. The eigenenergies $E_{\nu}(j,k_{z})$ of the $j=-1$ and 00 sectors are obtained analytically as
@@
The full Landau level spectrum (including all $j$) as a function of $k_{z}$ is shown in Fig. 3(a), where the lower panel presents a magnified plot of the flat-band region. The resulting spectrum can be classified into three main branches, labeled by $\nu=-1,0,1$, where $\nu=\pm 1$ correspond to flat bands, while $\nu=0$ corresponds to a dispersive branch. We find that the flat bands spread over finite energy range, $-\hbar\omega_{B}/4<E<\hbar\omega_{B}/4$. In the limit $k_{z}\rightarrow\infty$, all the $\nu=\pm 1$ levels converge to ±$\pm\hbar\omega_{B}/4$. We note that the maximum total angular momentum is estimated as $j_{\mathrm{max}}\sim l_{B}^{2}/d^{2}$ ensuring conservation of the total state density (see Appendix C). In Fig. 3, we set $j_{\mathrm{max}}=100$, which corresponds to $B\sim 12$ T for a typical pyrochlore material with $d\sim 5$ [97]. As $B$ decreases, $j_{\mathrm{max}}$ increases, and the gap edge of $\nu=+1$ ($\nu=-1$) branch shifts downward (upward) toward zero.
@@
Landau level spreading in large $k_{z}$ can be understood by a perturbation theory with respect to $H_{\mathrm{Z}}$. At the first order in $H_{\mathrm{Z}}$, the degeneracy between the flat bands with $L_{z}=\pm 1$ is lifted, shifting their energy to $\mp\hbar\omega_{B}/4$. This explains the splitting of the flat-band Landau levels into two branches in the $k_{z}\to\infty$ limit, as seen in Fig. 3. In this regime, the $\nu=\pm 1$ and 00 branches in Fig. 3 correspond to angular momentum $L_{z}=\mp 1$ and 00, respectively. As $k_{z}$ decreases, each branch of the flat-band Landau levels spreads. Using the second-order perturbation theory, the asymptotic behavior in the limit $j\ll k_{z}l_{B}$ can be written as
$\begin{split}E_{\pm 1}(j,k_{z})\approx\pm\frac{\hbar\omega_{B}}{4}\left(1-\frac{j+1/2\pm 1/2}{k_{z}^{2}l_{B}^{2}}\right)+\mathcal{O}\left(\frac{j}{k_{z}^{4}l_{B}^{4}}\right),\\ \end{split}$ (39)
@@
As approaching $k_{z}=0$, $\nu=+1$ levels rapidly contract to $E=0$, while the spread of $\nu=-1$ remains. At $k_{z}=0$, the Hamiltonian Eq. (30) is separated to block-diagonal matrices for the $L_{z}=\pm 1$ and $L_{z}=0$. The $L_{z}=\pm 1$ sector is equivalent to the continuum model for the kagome lattice hosting a 2D singular flat band and a dispersive band [4; 54], and the spread of the $\nu=-1$ Landau levels is equivalent to the LL spreading of the 2D singular flat band [4; 54]. The $L_{z}=0$ sector corresponds to a 2D trivial flat band with no singularity, consistent with the fully degenerate levels in $\nu=+1$ branch. As $k_{z}$ is continuously varied from $k_{z}=\infty$ to 00, the flat bands transition between different $L_{z}$ sectors: they are associated with $L_{z}=\pm 1$ in the $k_{z}\to\infty$ limit, whereas at $k_{z}=0$, they involve $L_{z}=\pm 1$ and 00.
Figure 4: Weyl semimetallic band structure associated with the orbital Zeeman effect. (a) Energy spectrum of $H_{\mathrm{eff+Z}}=H_{\mathrm{eff}}+H_{\mathrm{Z}}$. (b) Landau levels for $-1\leq j\leq 10$, overlaid on the spectrum in (a). (c) Schematic Landau level structure of a typical Weyl semimetal with the finite group velocity.

### v2

@@ I Introduction
Electronic systems featuring flat bands, where the bandwidth is vanishingly small compared to other relevant energy scales, have long been of interest in condensed matter physics, driven by their unusual single-particle properties [1, 2, 3, 4, 5], strongly correlated phases [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], and experimental realizations [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. A flat band can arise trivially when interatomic hopping is suppressed, leading to a so-called trivial flat band. In contrast, nontrivial flat bands emerge despite finite hopping amplitudes, as a result of destructive interference encoded in the lattice geometry. Prototypical examples include the dice [1], Lieb [7], and kagome [9] lattices. In such systems, the flat band is generally not isolated but instead touches a dispersive band at discrete points in momentum space. Considerable theoretical effort has been devoted to elucidating the mathematical origins and topological characteristics of these nontrivial flat bands [1, 8, 9, 10, 6, 46, 47, 2, 48, 49, 50, 51, 52, 53].
Recent studies have highlighted the role of nontrivial topology and quantum geometry in flat-band systems [54, 2, 55, 4, 56, 57, 58, 59, 60, 61, 60, 62], particularly in two-dimensional (2D) settings. Among these, singular flat bands [4, 54, 5, 55] have attracted growing attention. These are flat bands that remain exactly flat while touching a dispersive band at a singular point that cannot be removed without breaking the flatness. This singularity is protected by a discontinuity in the Bloch eigenvectors and is not merely a mathematical feature; it also has direct physical consequences. A striking effect appears in the Landau level structure under an external magnetic field. In a trivial flat band, the electronic states remain degenerate in the presence of a magnetic field, as the localized electrons are insensitive to the Aharonov–Bohm phase. By contrast, in a singular flat band, the magnetic field induces Landau levels spreading over a finite energy range reflecting the underlying geometric singularity [4, 54].
@@
A natural question is whether singular flat bands can exist in three dimensions (3D) and, if so, how they respond to an external magnetic field. The existence of 3D singular flat bands was first demonstrated as an extension of 2D studies [55]. A notable candidate is the pyrochlore lattice model, which hosts flat bands extending along all three momentum directions $(k_{x},k_{y},k_{z})$, with a band touching point at $\bm{k}=0$ [9, 2, 3]. This structure can be regarded as a 3D analog of the kagome lattice. Desipite its potential significance, the quantum geometric properties of these 3D flat bands remain largely unexplored.
In this work, we investigate the geometric singularity of the 3D flat band in the pyrochlore lattice and examine its response to an external magnetic field. Our analysis is based on a three-orbital effective continuum model, which captures the electronic structure near the threefold-degenerate band-touching point. We show that the touching point is a point-like topological singularity in three-dimensional momentum space, as illustrated in Fig. 1(a). At any $k$-point, the eigenvectors can be represented by three-component real vectors. Notably, the flat band in the 3D system is twofold degenerate unlike in 2D cases, and hence its eigenvectors span a two-dimensional subspace within the three-component vector space, illustrated as green flat planes in Fig. 1(a). The presence of the point-like singularity is manifested in a hedgehog structure of a normal director perpendicular to the plane of the flat-band manifold. Additionally, the quantum geometric tensor for the flat bands acquires a non-Abelian structure [63, 64, 65]: each of its component is a $2\times 2$ matrix, reflecting the intrinsic degeneracy of the flat bands.
@@
The three-dimensionality and intrinsic degeneracy introduce new features in the Landau level structure of the flat bands that are absent in 2D systems. For sufficiently large $|k_{z}|$, the degenerate flat bands are split into two branches by the orbital Zeeman effect, and the Landau levels within each branch remain nearly degenerate without energy spreading [Fig. 1(c)]. As $k_{z}$ decreases, the Landau levels in each branch acquire a finite energy spread, consistent with the increasing quantum metric. Near $k_{z}\sim 0$, hybridization between the two flat bands and the adjacent dispersive band gives rise to an intricate Landau level structure, potentially reflecting non-Abelian quantum geometry. We show that the full $k_{z}$-dependent behavior is well captured by an effective band model that incorporates the orbital Zeeman effect while neglecting full Landau quantization. Within this approximation, the system effectively behaves as a Weyl semimetal, and the Landau level spreading can be interpreted as a dispersion of the flat bands induced by orbital magnetism. The model also explains the presence of a chiral Landau level connecting the flat bands to the dispersive band — a characteristic feature of Weyl semimetals [66, 67, 68, 69, 70, 71].
@@
Lastly, while our effective continuum model for the pyrochlore lattice is constructed from three orbitals correponding to orbital angular momentum $\ell=1$, the framework can be naturally extended to arbtirary angular momentum quantum numbers $\ell$, described by $2\ell+1$ dimensional matrix. In particular, $\ell=3/2$ corresponds to Luttinger-Kohn-type model with parameters tuned to yield a flat band. While the original Luttinger-Kohn model with $\ell=3/2$ describes dispersive bands and has been studied in the context of various topological materials [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96], our work focuses on a flat band variant that can be generalized to arbitrary $\ell$. The extension demonstrates that the essential features of the 3D singular flat band identified in pyrochlore lattice, such as Landau level spreading and emergent Weyl semimetal like behevior, are robust across a wide class of band models.
@@ II.1 Effective continuum model
[…]
@@
$(H_{\bm{k}})_{X,X^{\prime}}$ $:=\langle\bm{k},\bm{\tau}_{X}|H|\bm{k},\bm{\tau}_{X^{\prime}}\rangle$
@@
$L_{x}=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&-i\\ 0&i&0\\ \end{array}\right),\ L_{y}=\left(\begin{array}[]{ccc}0&0&i\\ 0&0&0\\ -i&0&0\\ \end{array}\right),$ (14)
$L_{z}=\left(\begin{array}[]{ccc}0&-i&0\\ i&0&0\\ 0&0&0\\ \end{array}\right),$ (18)
which act as generators of the rotation operators in Eq. (7). These $L_{i}$ satisfy the standard algebra $L_{\alpha},L_{\beta}]=i\epsilon_{\alpha\beta\gamma}L_{\gamma}$, and represent effective orbital angular momentum associated with the sublattice degrees of freedom. Using them, the Hamiltonian in Eq. (6) can be rewritten as
@@ III.1 Landau level structure
We investigate the electronic structure of the 3D flat band system described by Eq. (19) under a uniform magnetic field $\bm{B}$ applied along the $z$-direction. We define the vector potential $\bm{A}$ satisfying $\bm{B}=\bm{\nabla}\times\bm{A}$. The Hamiltonian is then obtained by a standard process replacing the momentum components $k_{\alpha}$ in Eq. (19) with $\pi_{\alpha}/\hbar$ ($\alpha=x,y$), where $\bm{\pi}=-i\hbar\bm{\nabla}+e\bm{A}$ and the quadratic term is replaced with the symmetrized product $k_{\alpha}k_{\beta}\rightarrow(\pi_{\alpha}\pi_{\beta}+\pi_{\beta}\pi_{\alpha})/(2\hbar^{2})$. The momentum component along the magnetic field, $k_{z}$, remains a quantum number. We define the ladder operators $a=(\pi_{x}-i\pi_{y})/\sqrt{2\hbar eB}$ and $a^{\dagger}=(\pi_{x}+i\pi_{y})/\sqrt{2\hbar eB}$, which satisfy the commutation relation $a,a^{\dagger}]=1$, noting that $\pi_{x},\pi_{y}]=i\hbar eB$. As a result, the Hamiltonian in the magnetic field is written as
@@
Here, $\phi_{n}(x,y)$ denotes the wave function of $n$-th Landau levels in two-dimensional system with $n=0,1,2,\cdots$, where the index for the Landau level degeneracy in each $n$ is omitted. The $\vec{e}_{m}$ is a three-dimensional eigenvector of the orbital angular momentum $L_{z}$ with the eigenvalue m=1,0,1m=-1,0,1. The basis of Eq. (26) satisfies the following equations,
@@
For j=1j=-1 and 0, the eigenvectors reduce to one- and two-component forms, respectively, since the basis states $|k_{z},n,m\rangle$ with $n<0$ vanish. Correspondingly, the Hamiltonian matrices for j=1j=-1 and $j=0$ are given by the $1\times 1$ and $2\times 2$ lower-right submatrices of Eq. (30), respectively. The eigenenergies $E_{\nu}(j,k_{z})$ of the j=1j=-1 and 0 sectors are obtained analytically as
@@
The full Landau level spectrum (including all $j$) as a function of $k_{z}$ is shown in Fig. 3(a), where the lower panel presents a magnified plot of the flat-band region. The resulting spectrum can be classified into three main branches, labeled by $\nu=-1,0,1$, where $\nu=\pm 1$ correspond to flat bands, while $\nu=0$ corresponds to a dispersive branch. We find that the flat bands spread over finite energy range, $\hbar\omega_{B}/4<E<\hbar\omega_{B}/4$. In the limit $k_{z}\rightarrow\infty$, all the $\nu=\pm 1$ levels converge to $\pm\hbar\omega_{B}/4$. We note that the maximum total angular momentum is estimated as $j_{\mathrm{max}}\sim l_{B}^{2}/d^{2}$ ensuring conservation of the total state density (see Appendix C). In Fig. 3, we set $j_{\mathrm{max}}=100$, which corresponds to $B\sim 12$ T for a typical pyrochlore material with $d\sim 5$ Å [97]. As $B$ decreases, $j_{\mathrm{max}}$ increases, and the gap edge of $\nu=+1$ ($\nu=-1$) branch shifts downward (upward) toward zero.
@@
Landau level spreading in large $k_{z}$ can be understood by a perturbation theory with respect to $H_{\mathrm{Z}}$. At the first order in $H_{\mathrm{Z}}$, the degeneracy between the flat bands with $L_{z}=\pm 1$ is lifted, shifting their energy to $\mp\hbar\omega_{B}/4$. This explains the splitting of the flat-band Landau levels into two branches in the $k_{z}\to\infty$ limit, as seen in Fig. 3. In this regime, the $\nu=\pm 1$ and 0 branches in Fig. 3 correspond to angular momentum $L_{z}=\mp 1$ and 0, respectively. As $k_{z}$ decreases, each branch of the flat-band Landau levels spreads. Using the second-order perturbation theory, the asymptotic behavior in the limit $j\ll k_{z}l_{B}$ can be written as
[…]
@@
As approaching $k_{z}=0$, $\nu=+1$ levels rapidly contract to $E=0$, while the spread of $\nu=-1$ remains. At $k_{z}=0$, the Hamiltonian Eq. (30) is separated to block-diagonal matrices for the $L_{z}=\pm 1$ and $L_{z}=0$. The $L_{z}=\pm 1$ sector is equivalent to the continuum model for the kagome lattice hosting a 2D singular flat band and a dispersive band [4, 54], and the spread of the $\nu=-1$ Landau levels is equivalent to the LL spreading of the 2D singular flat band [4, 54]. The $L_{z}=0$ sector corresponds to a 2D trivial flat band with no singularity, consistent with the fully degenerate levels in $\nu=+1$ branch. As $k_{z}$ is continuously varied from $k_{z}=\infty$ to 0, the flat bands transition between different $L_{z}$ sectors: they are associated with $L_{z}=\pm 1$ in the $k_{z}\to\infty$ limit, whereas at $k_{z}=0$, they involve $L_{z}=\pm 1$ and 0.
Figure 4: Weyl semimetallic band structure associated with the orbital Zeeman effect. (a) Energy spectrum of $H_{\mathrm{eff+Z}}=H_{\mathrm{eff}}+H_{\mathrm{Z}}$. (b) Landau levels for $1\leq j\leq 10$, overlaid on the spectrum in (a). (c) Schematic Landau level structure of a typical Weyl semimetal with the finite group velocity.